OpenStudy (anonymous):

find the integral of : cosx/1-cos^2 dx

OpenStudy (anonymous):

$$\int \frac {\cos x} {1 - \cos^2 x} \, \mathrm{d}x$$ From the normalization identity, we know that \(\sin^2 x = 1 - \cos^2 x\).

OpenStudy (anonymous):

Thus: $$\int \frac{\cos x}{\sin^2 x} \, \mathrm{d}x$$

OpenStudy (anonymous):

Let us substitute \(u = \sin x\). Thus, \(\mathrm{d}u = \cos x \, \mathrm{d}x\).

OpenStudy (anonymous):

Sorry about previous answers, I meant: $$\int \frac{\mathrm{d}u}{u^2} = -\frac 1 2 u^{-1}$$

OpenStudy (anonymous):

Okay, let's substitute \(u = \sin x\).

OpenStudy (anonymous):

The answer is $$-\frac{1}{2 \sin x}$$

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

another one please

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

integral sinx(2cscx+sinx)dx

OpenStudy (anonymous):

$$\int (\sin x)(2 \csc x + \sin x) \, \mathrm{d}x$$

OpenStudy (anonymous):

\(\csc x = \frac 1 {\sin x}\).

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1+sin^2x

OpenStudy (anonymous):

Kind of. You forgot the 2. $$\int 2 + \sin^2 x \, \mathrm{d}x$$.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the last one?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Have you got the answer to the previous integral?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x+...?

OpenStudy (anonymous):

?????????????

OpenStudy (raden):